OpenStudy (anonymous):

pam has 120 pesos and is spending 5 pesos every week. lorenzo has 20 pesos and is saving 7.50 pesos every week, when they have the same amount of money? (answer it using system of linear equations) please solve it. i need it tom. thanks

OpenStudy (anonymous):

p = 120 - 5 w p = 20 + 7.50 w With p (pesos of both people), w number of week. We have to find w,so 120 - 5 w = 20 + 7.50 w Can you solve it from there ?

OpenStudy (anonymous):

no. because i cant understand it. please answer it by using system of linear equations.thank you

OpenStudy (anonymous):

\[ \left\{\begin{matrix} p = 120 - 5 w\\ p = 20 + 7.50 w \end{matrix}\right. \] Here is your system of linear equation. 120 - 5 w = 20 + 7.5 w 120 - 20 = 7.5 w + 5 w 100 = 12.5 w w = 100/12.5 w = 8

OpenStudy (anonymous):

So both will have the same money after 8 weeks